# On the Hurwitz Stability of Hurwitz-Type Matrix Polynomials

### v1

@@ 2.3 Stieltjes orthogonal matrix polynomials
$s^{(2m)}_{\scriptscriptstyle{\left(\begin{smallmatrix}0\\ I\end{smallmatrix}\right)}}(z):=-\frac{Q_{1,m}^{*}(\bar{z})}{P_{1,m}^{*}(\bar{z})}.$ (58)
By using [23] and an equivalent form of the resolvent matrix $U^{(2m)}$, it is proven that
$s^{(2m)}_{\scriptscriptstyle{\left(\begin{smallmatrix}I\\ 0\end{smallmatrix}\right)}}(x)<s^{(2m)}_{\scriptscriptstyle{\left(\begin{smallmatrix}0\\ I\end{smallmatrix}\right)}}(x),\quad x\in(-\infty,0).$ (59)
In [7, Theorem 3.4], a representation of extremal solutions (58) and (57) in terms of matrix continued fractions was given. We reproduce this result in the following remark.
Remark 2.11.
Let $P_{k,j}$, $Q_{k,j}$, for $k=1,2$ be as in Definition 2.8 and for $M_{k}$, $L_{k}$ as in (35) through (37). The extremal solutions to the TSMM problem defined in $z\in\mathbb{C}\setminus[0,\infty)$ then take the following form:
$s^{(2m)}_{\scriptscriptstyle{\left(\begin{smallmatrix}I\\ 0\end{smallmatrix}\right)}}(z)=-\frac{Q_{2,m}^{*}(\bar{z})}{zP_{2,m}^{*}(\bar{z})}=$ $\cfrac{I_{q}}{-zM_{0}+\cfrac{I_{q}}{L_{0}+\cfrac{I_{q}}{\ddots\,-zM_{m-1}+\cfrac{I_{q}}{L_{m-1}-z^{-1}M_{m}^{-1}}}}},$ (60)
$s^{(2m)}_{\scriptscriptstyle{\left(\begin{smallmatrix}0\\ I\end{smallmatrix}\right)}}(z)=-\frac{Q_{1,m}^{*}(\bar{z})}{P_{1,m}^{*}(\bar{z})}=$ $\cfrac{I_{q}}{-zM_{0}+\cfrac{I}{L_{0}+\cfrac{I_{q}}{\ddots+L_{m-2}+\cfrac{I_{q}}{-zM_{m-1}+L_{m-1}^{-1}}}}}.$ (61)
2.4 Relatively prime matrix polynomials
Recall representation (2) of an HTM polynomial ${\bf f}_{n}$. To explain that the pairs $({\bf h}_{2m}(z),{\bf g}_{2m}(z))$ (resp. $({\bf h}_{2m+1}(z),z{\bf g}_{2m+1}(z))$) are coprime, as in Theorem 3.5, we reproduce Definition 1.15.7 and Theorem 1.15.5 of [44] related to relatively right prime or right coprime $q\times q$ matrix polynomials.
Definition 2.12.
a) A $q\times q$ matrix $W(z)$ is called unimodular if $\det W(z)$ is equal to a nonzero constant quantity [45, page 370], independent of $z$.
b) Two $q\times q$ polynomial matrices $A(z)$ and $B(z)$ are called left (right) or row (column) equivalent if and only if one can be obtained from the other as a result of a finite number of elementary operations performed on its rows (columns). Specifically, $B(z)=L(z)A(z)$ (resp. $B(z)=A(z)P(z)$) where $L(z)$ (resp. $P(z)$) is the product of unimodular matrices corresponding to elementary operations on rows (columns). See [44, Definition 1.15.7].
Definition 2.13.
A matrix $L(z)\in\mathbb{C}^{q\times q}$ is called a left common divisor of $q\times q$ matrices $A(z)$ and $B(z)$ if and only if there exist matrices $A_{1}(z),B_{1}(z)\in\mathbb{C}^{q\times q}$ such that
$A(z)=L(z)A_{1}(z)\qquad\mbox{and}\qquad B(z)=L(z)B_{1}(z).$ (62)
A matrix $P(z)\in\mathbb{C}^{q\times q}$ is called a right common divisor of $q\times q$ matrices $A(z)$ and $B(z)$ if and only if there exist matrices $A_{2}(z),B_{2}(z)\in\mathbb{C}^{q\times q}$ such that
$A(z)=A_{2}(z)P(z)\qquad\mbox{and}\qquad B(z)=B_{1}(z)P(z).$ (63)
Definition 2.14.
Let $q\times q$ matrix-valued $A(z)$ and $B(z)$ are called relatively left prime (resp. right prime) if and only if only their left (resp. right) common divisors are unimodular matrices.
The following theorem appears in [44, Theorem 1.15.5], [61, Theorem 6.1] and [66, Theorem 2.5.16].
Theorem 2.15.
[…]
Matrix polynomials $A(z)$, $B(z)\in\mathbb{C}^{q\times q}$ are right coprime if and only if $\begin{pmatrix}A(z)\\ B(z)\end{pmatrix}$ and $\begin{pmatrix}I_{q}\\ 0_{q}\end{pmatrix}$ are left equivalent.
Related to the coprimeness of two matrix polynomials, we reproduce [45, Lemma 6.3-5], [3, page 585], [52, page 26].
Lemma 2.16.
Matrices $A(z)$ and $B(z)$ are right coprime if and only if polynomial matrices $X(z)$ and $Y(z)$ exist such that
$X(z)A(z)+Y(z)B(z)=I.$ (64)
Equality (64) is called Bézout’s identity.
Identity (64) also appears in [28, page 21] and [65, Corollary 4.1.4] in the context of control systems theory.
@@ 2.5 The Bezoutian and inertia for matrix polynomials
In this subsection, we reproduce some notions related to the inertia for a matrix polynomial [53, 54]. The inertia for scalar polynomials was considered in [51, page 462].
@@
is called inertia with respect to the real axis, where $\gamma_{+}(L)$ (resp. $\gamma_{-}(L)$) is the number of zeros of $\det(L(\lambda))$ in the open upper (resp. open lower) half-plane, and $\gamma_{0}(L)$ is the number of zeros of $\det(L(\lambda))$ in the real axis. Now we reproduce [53, Corollary 2.2].
@@
The generalized Bezoutian form associated with the quadruple ${A,B,C,D}$ is defined as
$\Gamma(x,y)=\frac{1}{x-y}\left(A(x)D(y)-B(x)G(y)\right).$ (71)
@@
where $s_{{\bf f}_{n}}(z)$ denotes the left-hand side of (10), (11). See Remark 3.1.
@@
3 Properties of the Hurwitz-type matrix polynomials
In this section, we consider properties of Hurwitz-type matrix polynomials.
Remark 3.1.
Let $s_{{\bf f}_{n}}(z)$ denote the left-hand side of (10), (11). By taking into account that ${\bf c}_{j}$ and ${\bf d}_{j}$ are positive definite, equality (72) is valid.
Theorem 3.2.
Let $n$ be a natural number such that $n\geq 2$. Furthermore, let ${\bf f}_{n}$ be a matrix-valued polynomial and ${\bf h}_{n}$ (resp. ${\bf g}_{n}$) be the ${\bf h}_{n}$ (resp. ${\bf g}_{n}$) parts of ${\bf f}_{n}$. The assertions below are valid.
  • (a)
    The HTM polynomial ${\bf f}_{2m}$ (resp. ${\bf f}_{2m+1}$) is associated with the finite sequence of Markov parameters or moments $(s_{j})_{j=0}^{2m-1}$ (resp. $(s_{j})_{j=0}^{2m}$), which are Hermitian matrices.
  • (b)
    Let $n=2m$ (resp. $n=2m+1$) with $m\geq 1$, and let $H_{1,m-1}$ and $H_{2,m-1}$ (resp. $H_{1,m}$ and $H_{2,m-1}$) be constructed via Hermitian Markov parameters of the polynomial ${\bf f}_{n}$. The matrix-valued polynomial ${\bf f}_{n}$ is an HTM polynomial if and only if the block Hankel matrices $H_{1,m-1}$ and $H_{2,m-1}$ (resp. $H_{1,m}$ and $H_{2,m-1}$) are positive definite matrices.
  • (c)
    Let $n=2m$ (resp. $n=2m+1$) with $m\geq 1$, and let $\widehat{H}_{1,m-1}$ and $\widehat{H}_{2,m-1}$ (resp. $\widehat{H}_{1,m}$ and $\widehat{H}_{2,m-1}$) be the Schur complements of Hermitian block Hankel matrices $H_{1,m-1}$ and $H_{2,m-1}$ (resp. $H_{1,m}$ and $H_{2,m-1}$) , respectively. Thus, ${\bf f}_{n}$ is an HTM polynomial if and only if the Schur complements $\widehat{H}_{1,m-1}$ and $\widehat{H}_{2,m-1}$ (resp. $\widehat{H}_{1,m}$ and $\widehat{H}_{2,m-1}$) are positive definite matrices. The matrices $\widehat{H}_{1,m-1}$ and $\widehat{H}_{2,m-1}$ (resp. $\widehat{H}_{1,m}$ and $\widehat{H}_{2,m-1}$) are constructed via the Markov parameters of the polynomial ${\bf f}_{n}$.
  • (d)
    Let $n=2m$ (resp. $n=2m+1$) with $m\geq 1$. The matrix-valued polynomial ${\bf f}_{n}$ is an HTM polynomial if and only if the rational matrix-valued polynomial $s_{{\bf f}_{n}}(-z)$, as in (72), is an extremal solution to the TSMM problem, as in (58) (resp. (57)).
  • (e)
    Let $n=2m$ (resp. $n=2m+1$). The pair of sequences $(({\bf c}_{k})_{k=0}^{m},({\bf d}_{k})_{k=0}^{m-1})$ $(resp.(({\bf c}_{k})_{k=0}^{m},({\bf d}_{k})_{k=0}^{m}))$ as in Definition 1.1 A and A the pair of sequences $((M_{k})_{k=0}^{m},(L_{k})_{k=0}^{m-1})$ $(resp.((M_{k})_{k=0}^{m},(L_{k})_{k=0}^{m}))$ as in (35)-(37) coincide.
  • (f)
    For $n\geq 2$, the matrix polynomial ${\bf f}_{n}$ is an HTM polynomial if and only if ${\bf f}_{n}$ has the following representation:
    where $P_{k,m}$ and $Q_{k,m}$ for $k=1,2$ are polynomials defined in Definition 2.8. The matrix polynomials $P_{k,m}$ and $Q_{k,m}$ are constructed via the Markov parameters of the polynomial ${\bf f}_{n}$.
Proof.
Part (a) is proven in [8, Theorem 7.9]. Part (b) is proven in [8, Theorem 7.10]. Part (c) is established using part (b) and [36, Corollary 1], which demonstrate that the positive definiteness of the block matrix $H_{k,j}$ is equivalent to that of its Schur complement $\widehat{H}_{k,j}$. Part (d) follows from [7, Theorem 7.12] and the uniqueness of the extremal solutions (61) and (60). Part (e) is proven in [8, Theorem 7.12]. The proof of part (f) follows directly from (2), part (b) and the equalities
${\bf h}_{n}(z)=$ $(-1)^{m}Q_{2,m}^{*}(-\bar{z})\quad\mbox{and}\quad{\bf g}_{n}(z)=(-1)^{m}P_{2,m}^{*}(-\bar{z}),\qquad n=2m+1.$ (76)
Equalities (75) and (76) were proven in the proof of [7, Proposition 7.8]. ∎
The corresponding distribution functions of the Stieltjes transforms $s^{(2m)}_{\scriptscriptstyle{\left(\begin{smallmatrix}I\\ 0\end{smallmatrix}\right)}}(z)$ and $s^{(2m)}_{\scriptscriptstyle{\left(\begin{smallmatrix}0\\ I\end{smallmatrix}\right)}}(z)$, associated with the HTM polynomials as given in (60) and (61), can be represented by monotonically non-decreasing, piecewise constant matrix function [4].
A $q\times q$ matrix $\widetilde{\sigma}(\cdot)$ defined on $[0,\infty)$ is called monotonically nondecreasing if $\widetilde{\sigma}(t)$ is self-adjoint on $[0,\infty)$ and $\widetilde{\sigma}(t_{2})-\widetilde{\sigma}(t_{1})\geq 0$ for $t_{1}\leq t_{2}$. For $n=2m+1$ (resp. $n=2m$), the points of discontinuity of the distribution function $\widetilde{\sigma}$ that correspond to the measure $\sigma$ are given by the roots of polynomial $\det(zP_{2,m}^{*}(-\bar{z}))$ (resp. $\det(P_{1,m}^{*}(-\bar{z}))$). On the other hand, polynomials (75) and (76), which are constructed using the moments $s_{j}$, can be obtained from an absolutely continuous distribution on $[0,\infty)$ that corresponds to a positive measure on $[0,\infty)$, as shown in the next example.
Example 3.3.
Let $\widetilde{\sigma}(x)=\begin{pmatrix}4-2e^{-x/2}&ie^{-x}\\ -ie^{-x}&2-e^{-x}\\ \end{pmatrix}$ be defined on $[0,\infty)$. We construct the moments $s_{0}=\begin{pmatrix}2&-i\\ i&1\end{pmatrix}$, $s_{1}=\begin{pmatrix}4&-i\\ i&1\end{pmatrix}$, $s_{2}=\begin{pmatrix}16&-2i\\ 2i&2\end{pmatrix},$ $s_{3}=\begin{pmatrix}96&-6i\\ 6i&6\end{pmatrix}$, $s_{4}=\begin{pmatrix}768&-24i\\ 24i&24\end{pmatrix}$ of $\widetilde{\sigma}(x)$. The sequence $(s_{j})_{j=0}^{4}$ is a Stieltjes positive definite sequence. By using these moments and Definition 2.8, we derive $P_{2,2}$ and $Q_{2,2}$. From (74), we obtain the HMT polynomial
${\bf f}_{5}(z)=z^{5}I_{2}+A_{1}z^{4}+A_{2}z^{3}+A_{3}z^{2}+A_{4}z+A_{5},$ (77)
$A_{4}=\begin{pmatrix}\frac{1158}{37}&0\\ -\frac{936}{37}i&6\\ \end{pmatrix}$ and $A_{5}=\begin{pmatrix}\frac{292}{37}&-2i\\ 2i&2\\ \end{pmatrix}.$ These coefficients $A_{j}$ clearly satisfy Condiction C (14). Therefore, by Theorem 5.3, the polynomial ${\bf f}_{5}$ is a Hurwitz polynomial.
3.1 Coprimeness of pairs $({\bf h}_{n}(z),g_{n}(z))$ and $({\bf h}_{n}(z),zg_{n}(z))$
We paraphrase [7, Proposition 5.1 (a)] to apply it in Theorem 3.5.
Proposition 3.4.
Let $(s_{j})_{j=0}^{2m-1}$ (resp. $(s_{j})_{j=0}^{2m}$) be Stieltjes positive sequences. Furthermore, let $P_{k,j}$ and $Q_{k,j}$ be as in Definition 2.8. Moreover, let $\widehat{H}_{1,j}$ be as in (33). Thus, the following equalities hold:
$P_{1,j}(z)Q_{2,j}^{*}(\bar{z})-zQ_{1,j}(z)P_{2,j}^{*}(\bar{z})=$ $\widehat{H}_{1,j},$ (78)
$P_{k,j}(z)Q_{2,j}^{*}(\bar{z})-Q_{k,j}(z)P_{k,j}^{*}(\bar{z})=$ $0_{q}.$ (79)
In the following result, we consider the coprimeness of the ${\bf h}_{n}$ and ${\bf g}_{n}$ parts of a given polynomial ${\bf f}_{n}$. To prove this, we utilize Bézout’s identity (64), which involves orthogonal matrix polynomials $P_{k,j}$ on $[0,+\infty)$ and their second-kind polynomials $Q_{k,j}$.
Theorem 3.5.
(a) Let ${\bf f}_{2m}(z)$ be an HTM polynomial. Thus, its corresponding ${\bf h}_{2m}(z)$ and ${\bf g}_{2m}(z)$ are right coprime polynomials.
(b) Let ${\bf f}_{2m+1}(z)$ be an HTM polynomial. Thus, its corresponding ${\bf h}_{2m+1}(z)$ and $z{\bf g}_{2m+1}(z)$ are right coprime polynomials.
Proof.
From Theorem 3.2(a), we deduce that the polynomial ${\bf f}_{2m+1}$ (resp. ${\bf f}_{2m}$) is associated with the $(s_{j})_{j=0}^{2m}$ (resp $(s_{j})_{j=0}^{2m-1}$) Markov parameters. For the polynomial ${\bf f}_{2m}$, we apply Proposition 2.5 and find a parameter $s_{2m}$ that satisfies Inequality (21). We exploit the fact that $(s_{j})_{j=0}^{2m-1}$ is a Hankel nonnegative definite extendable sequence and that $s_{2m}$ represents the indicated extension. See Definition 2.3. Consequently, polynomials $P_{1,m}(z)$ and $Q_{2,m}^{*}(\bar{z})$ are well defined for ${\bf f}_{2m}$ and ${\bf f}_{2m+1}$. Let $n=2m+1$.
By using direct calculations and (76), we rewrite (78) for $j=m$ as
$X_{1}(z){\bf h}_{2m+1}(z)+Y_{1}(z)(z\,{\bf g}_{2m+1}(z))=I,$ (80)
where $X_{1}(z):=(-1)^{m}\widehat{H}_{1,m}^{-1}P_{1,m}(-z)$ and $Y_{1}(z):=(-1)^{m}\widehat{H}_{1,m}^{-1}Q_{1,m}(-z)$. Recall that the Schur complement $\widehat{H}_{1,m}$ is invertible because $H_{1,m}$ is positive definite.
For $n=2m$, by taking the adjoint of (78) and by replacing $z$ with $\bar{z}$, we obtain
$X_{2}(z){\bf h}_{2m}(z)+Y_{2}(z){\bf g}_{2m}(z)=I,$ (81)
where $X_{2}(z):=(-1)^{m}\widehat{H}_{1,m}^{-1}Q_{2,m}(-z)$ and $Y_{2}(z):=(-1)^{m}z\widehat{H}_{1,m}^{-1}P_{2,m}(-z)$.
From (80) (resp. (81) ) and Lemma 2.16, we observe that the polynomials ${\bf h}_{2m+1}(z)$ and $z\,{\bf g}_{2m+1}(z)$ (resp. ${\bf h}_{2m}(z)$ and ${\bf g}_{2m}(z)$) are right coprime. ∎
Now we present an alternative proof of Theorem 3.5. Although different regarding the inverse of the matrix $U^{(2m)}$ as in (46), a similar proof was provided in [12].
Alternative proof of Theorem 3.5. The inverse of the resolvent matrix (46) [7] is the following:
$U^{{(2m)}^{-1}}(z)=\begin{pmatrix}P_{1,m}^{-1}(0)P_{1,m}(z)&P_{1,m}^{-1}(0)Q_{1,m}(z)\\ zQ_{2,m}^{-1}(0)P_{2,m}(z)&Q_{2,m}^{-1}(0)Q_{2,m}(z)\end{pmatrix},\quad z\in\mathbb{C}.$ (82)
To prove (a), we introduce the matrix
$X_{3}(z):={\rm diag}(\widehat{H}_{1,m}^{-1}P_{1,m}(0),I_{q})\,U^{{(2m)}^{-1}}(z)\,{\rm diag}((-1)^{m}I_{q},(-1)^{m+1}I_{q}).$ (83)
With (78), (79) and (76), it can be readily verified that
$X_{3}(z)\begin{pmatrix}{\bf h}_{2m+1}(z)\\ z\,{\bf g}_{2m+1}(z)\end{pmatrix}=\begin{pmatrix}I_{q}\\ 0\end{pmatrix}.$ (84)
By (55) and Definition 2.12 a), the $2q\times 2q$ matrix-valued polynomial $X_{3}(z)$ is unimodular. The fact that the pair $({\bf h}_{2m+1}(z),z{\bf g}_{2m+1}(z))$ is left coprime follows from Theorem 2.15, and part (a) is proved.
Now we prove part (b). Let $K_{1}:=\begin{pmatrix}0&\widehat{H}_{1,m}^{-1}Q_{2,m}(0)\\ I&0\end{pmatrix}$. Furthermore, let
$X_{4}(z):=K_{1}U^{{(2m)}^{-1}}(z)\,{\rm diag}((-1)^{m}I,(-1)^{m}I).$ (85)
By incorporating (78), (79) and (75), it can be readily verified that
$X_{4}(z)\begin{pmatrix}{\bf h}_{2m}(z)\\ {\bf g}_{2m}(z)\end{pmatrix}=\begin{pmatrix}I_{q}\\ 0_{q}\end{pmatrix}.$ (86)
By offering similar arguments as in the proof of part (a), it follows that the pair $({\bf h}_{2m}(z),{\bf g}_{2m}(z))$ is left coprime. Thus, part (b) is proven.
4 Bezoutian and HTM polynomials
In this section, under the commutative-like Condition C (14), we derive an explicit form of the Bezoutian for the quadruple (${\bf f}_{n}^{*}(\bar{x}),{\bf f}_{n}(-x),{\bf f}_{n}^{*}(-\bar{x}),{\bf f}_{n}(x)$). Instead of working directly with the Bezoutian form
@@ 4 Bezoutian and HTM polynomials
associated with the quadruples $(x{\bf g}_{n}^{*},{\bf h}_{n},{\bf h}_{n}^{*},x{\bf g}_{n})$ and $({\bf g}_{n}^{*},{\bf h}_{n},{\bf h}_{n}^{*},{\bf g}_{n})$. Recall that a similar decomposition of the Bezoutian for scalar polynomials was previously studied; see [48, page 293] and [1, page 257]. Following [2] to compute the Bezoutian forms ${\mathscr{G}}_{n}^{(1)}(x,y)$ and ${\mathscr{G}}_{n}^{(2)}(x,y)$ we employ both representations (72) of the rational matrices (10) and (11).
@@
Let Condition C, given by (14), be satisfied. Furthermore, let ${\mathscr{F}}_{n}$, ${\mathscr{G}}_{n}^{(1)}$ and ${\mathscr{G}}_{n}^{(2)}$ be as defined in (87), (88) and (89), respectively. The functions ${\mathscr{G}}_{n}^{(1)}$, ${\mathscr{G}}_{n}^{(2)}$ and ${\mathscr{F}}_{n}$ are polynomials on $x$ and $y$.
@@
Let $n=2m$, and let
$\widetilde{V}_{[0,2m]}(x):=\left(\bar{x}^{2m}I_{q},\bar{x}^{2m-2}I_{q},\ldots,\bar{x}^{2}I_{q},I_{q}\right)^{\mathsf{*}}.$ (91)
For $i,j=0,2,4,\ldots,2m$, let
$\alpha_{ij}:=-A_{i}^{*}A_{j+1}+A_{i+1}^{*}A_{j},$ (92)
where $A_{j}$ is a coefficient of ${\bf f}_{n}$ and we set $A_{2m+1}=0$. Additionally, let
${\mathscr{A}}_{2m}:=\begin{pmatrix}0&\alpha_{0,2}&\alpha_{0,4}&\ldots&\alpha_{0,2m-2}&\alpha_{0,2m}\\ -\alpha_{0,2}&0&\alpha_{2,4}&\ldots&\alpha_{2,2m-2}&\alpha_{2,2m}\\ -\alpha_{0,4}&-\alpha_{2,4}&0&\ldots&\alpha_{4,2m-2}&\alpha_{4,2m}\\ \ldots&\ldots&\ldots&\ldots&\ldots&\ldots\\ -\alpha_{0,2m-2}&-\alpha_{2,2m-2}&-\alpha_{4,2m-2}&\ldots&0&\alpha_{2m-2,2m}\\ -\alpha_{0,2m}&-\alpha_{2,2m}&-\alpha_{4,2m}&\ldots&-\alpha_{2m-2,2m}&0\end{pmatrix}$ (93)
By using (5) , (9), (91) , (92), (93), as well as Condition C (14), we have
$x^{2}{\bf g}_{n}^{*}(\bar{x}^{2}){\bf h}_{n}(y^{2})-y^{2}{\bf h}_{n}^{*}(\bar{x}^{2}){\bf g}_{n}(y^{2})=\widetilde{V}_{[0,2m]}^{*}(\bar{x}){\mathscr{A}}_{2m}\widetilde{V}_{[0,2m]}(y)$
$=\alpha_{0,2m}(x^{2m}-y^{2m})+\alpha_{0,2m-2}x^{2}y^{2}(x^{2m-2}-y^{2m-2})+\ldots+\alpha_{2m-2,2m}(x^{2}-y^{2}).$
It follows that ${\mathscr{F}}_{n}$ is a polynomial in $x$ and $y$. From the last equality and right-hand side of (88), we see that ${\mathscr{G}}_{2m}^{(1)}(x,y)$ is polynomial on $x$ and $y$.
Similarly, it can be shown that ${\mathscr{G}}_{2m}^{(2)}$, ${\mathscr{G}}_{2m+1}^{(1)}$ and ${\mathscr{G}}_{2m+1}^{(2)}$ are polynomials in $x$ and $y$. From (90), it follows that ${\mathscr{F}}_{n}$ is a polynomial in $x$ and $y$. ∎
@@
Let Condition C, given by (14), be satisfied. Furthermore, let ${\mathscr{F}}_{n}$, ${\mathscr{G}}_{n}^{(1)}$ and ${\mathscr{G}}_{n}^{(2)}$ be as in (87), (88) and (89), respectively. Thus, for $n=2m$, we have
${\mathscr{G}}_{2m}^{(1)}(x,y)=$ ${\bf h}_{2m}^{*}(\bar{x}^{2})\left(x^{-2}I_{q},x^{-4}I_{q},\ldots\right)H^{(1)}_{\infty}\left(y^{-2}I_{q},y^{-4}I_{q},\ldots\right)^{\mathsf{T}}{\bf h}_{2m}(y^{2}),$ (95)
${\mathscr{G}}_{2m}^{(2)}(x,y)=$ $-{\bf h}_{2m}^{*}(\bar{x}^{2})\left(x^{-1}I_{q},x^{-3}I_{q},\ldots\right)H^{(0)}_{\infty}\left(y^{-1}I_{q},y^{-3}I_{q},\ldots\right)^{\mathsf{T}}{\bf h}_{2m}(y^{2}).$ (96)
@@
${\mathscr{G}}_{2m+1}^{(1)}(x,y)=$ ${\bf g}_{2m+1}^{*}(\bar{x}^{2})\left(I_{q},x^{-2}I_{q},\ldots\right)H^{(0)}_{\infty}\left(I_{q},y^{-2}I_{q},\ldots\right)^{\mathsf{T}}{\bf g}_{2m+1}(y^{2}),$ (97)
${\mathscr{G}}_{2m+1}^{(2)}(x,y)=$ $-{\bf g}_{2m+1}^{*}(\bar{x}^{2})\left(x^{-1}I_{q},x^{-3}I_{q},\ldots\right)H^{(1)}_{\infty}\left(y^{-1}I_{q},y^{-3}I_{q},\ldots\right)^{\mathsf{T}}$
$\cdot{\bf g}_{2m+1}(y^{2}),$ (98)
where the superscript $\mathsf{T}$ signifies the transpose of a matrix.

### v3

@@ 2.2 Stieltjes orthogonal matrix polynomials
${\bf h}_{n}(z)=$ $(-1)^{m}Q_{2,m}^{*}(-\bar{z})\quad\mbox{and}\quad{\bf g}_{n}(z)=(-1)^{m}P_{2,m}^{*}(-\bar{z}),\qquad n=2m+1$ (26)
which are established in the proof of [5, Proposition 7.8].
The notation below was introduced in [5, Lemma 2.6]. If $n=2m$, we set ${\mathscr{N}}_{{\bf h}_{n}}:=\{z\in{\mathbb{C}}:\,\det{\bf h}_{n}(z)=0\}$ , whereas for $n=2m+1$ we set ${\mathscr{N}}_{{\bf g}_{n}}:=\{z\in{\mathbb{C}}:\,\det{\bf g}_{n}(z)=0\}$. Furthermore, we define $\eta_{n}:=\max\{|z|:\,z\in{\mathscr{N}}_{{\bf h}_{n}}\}$, respectively, $\rho_{n}:=\max\{|z|:\,z\in{\mathscr{N}}_{{\bf g}_{n}}\}$. In the proof of [5, Theorem 6.1], the so–called extremal solutions to the TSMM problem are expressed as
$G_{\sigma_{2m,\max}}(z)=-\frac{Q_{1,j}^{*}(\bar{z})}{P_{1,j}^{*}(\bar{z})}=\frac{{\bf g}_{2m}(-z)}{{\bf h}_{2m}(-z)}$ (27)
and
$G_{\sigma_{2m,\min}}(z)=-\frac{Q_{2,j}^{*}(\bar{z})}{zP_{2,j}^{*}(\bar{z})}=\frac{{\bf h}_{2m+1}(-z)}{(-z){\bf g}_{2m+1}(-z)}.$ (28)
Note that the matrix function $G_{\sigma_{2m,\max}}$ (resp. $G_{\sigma_{2m,\min}}$) is defined for all $z\in{\mathbb{C}}\setminus[0,\infty)$ with $|z|>\eta_{2m}$ (resp. for all $z\in{\mathbb{C}}\setminus[0,\infty)$ with $|z|>\rho_{2m+1}$).
@@ 2.3 The Bezoutian and inertia for matrix polynomials
In this subsection, we reproduce some notions related to the inertia for a matrix polynomial [43, 44]. The inertia for scalar polynomials was considered in [42, page 462].
@@
is called inertia with respect to the real axis, where $\gamma_{+}(L)$ (resp. $\gamma_{-}(L)$) is the number of zeros of $\det(L(\lambda))$ in the open upper (resp. open lower) half-plane, and $\gamma_{0}(L)$ is the number of zeros of $\det(L(\lambda))$ in the real axis. Now we reproduce [43, Corollary 2.2].
@@
The generalized Bezoutian form associated with the quadruple A,B,C,D{A,B,C,D} is defined as
$\Gamma(x,y)=\frac{1}{x-y}\left(A(x)D(y)-B(x)G(y)\right).$
@@
where $s_{{\bf f}_{n}}(z)$ denotes the left-hand side of (9), (10).
@@
3 Bezoutian and Hurwitz-type matrix polynomials
In this section, using the so-called matrix symmetrizers (55) and block Hanckel matrices $H_{1,j}$ and $H_{2,j}$, we derive an explicit form of the Bezoutian for the quadruple (${\bf f}_{n}^{*}(\bar{x}),{\bf f}_{n}(-x),{\bf f}_{n}^{*}(-\bar{x}),{\bf f}_{n}(x)$).
Instead of working directly with the Bezoutian form
@@ 3 Bezoutian and Hurwitz-type matrix polynomials
associated with the quadruples $(x{\bf g}_{n}^{*},{\bf h}_{n},{\bf h}_{n}^{*},x{\bf g}_{n})$ and $({\bf g}_{n}^{*},{\bf h}_{n},{\bf h}_{n}^{*},{\bf g}_{n})$. Recall that a similar decomposition of the Bezoutian for scalar polynomials was previously studied; see [40, page 293] and [1, page 257]. Following [2] to compute the Bezoutian forms ${\mathscr{G}}_{n}^{(1)}(x,y)$ and ${\mathscr{G}}_{n}^{(2)}(x,y)$ we employ both representations (35) of the rational matrices (9) and (10).
@@
Let ${\mathscr{F}}_{n}$, ${\mathscr{G}}_{n}^{(1)}$ and ${\mathscr{G}}_{n}^{(2)}$ be as defined in (36), (37) and (38), respectively. The functions ${\mathscr{F}}_{n}$, ${\mathscr{G}}_{n}^{(1)}$ and ${\mathscr{G}}_{n}^{(2)}$ are polynomials on $x$ and $y$.
@@
Using (24) and (23), we have
${\bf f}_{n}(\bar{x}){\bf f}_{n}(-x)-{\bf f}_{n}(-\bar{x}){\bf f}_{n}(x)$
$=2x\left\{\begin{array}[]{lc}P_{1,m}(-x^{2})Q_{1,m}^{*}(-x^{2})-Q_{1,m}(-x^{2})P_{1,m}^{*}(-x^{2}),&n=2m\\ P_{2,m}(-x^{2})Q_{2,m}^{*}(-x^{2})-Q_{2,m}(-x^{2})P_{2,m}^{*}(-x^{2}),&n=2m+1\end{array}\right$.
$=0.$
By [15, Chapter 1], any polynomial vanishing on the diagonal $x=y$ is divisible by xyx-y. Since ${\bf f}_{n}^{*}(\bar{x}){\bf f}_{n}(-y)-{\bf f}_{n}^{*}(-\bar{x}){\bf f}_{n}(y)$ has this property, ${\mathscr{F}}_{n}(x,y)$ is a polynomial. Arguing similarly with (24), (23), and [15, Chapter 1], we show that ${\mathscr{G}}_{n}^{(1)}$ and ${\mathscr{G}}_{n}^{(2)}$ are polynomials in $x$ and $y$. ∎
@@
Let ${\mathscr{F}}_{n}$, ${\mathscr{G}}_{n}^{(1)}$ and ${\mathscr{G}}_{n}^{(2)}$ be as in (36), (37) and (38), respectively. Thus, for $n=2m$, we have
${\mathscr{G}}_{2m}^{(1)}(x,y)=$ ${\bf h}_{2m}^{*}(\bar{x}^{2})\left(x^{-2}I_{q},x^{-4}I_{q},\ldots\right)H^{(1)}_{\infty}\left(\bar{y}^{-2}I_{q},\bar{y}^{-4}I_{q},\ldots\right)^{*}{\bf h}_{2m}(y^{2}),$ (41)
${\mathscr{G}}_{2m}^{(2)}(x,y)=$ $-{\bf h}_{2m}^{*}(\bar{x}^{2})\left(x^{-1}I_{q},x^{-3}I_{q},\ldots\right)H^{(0)}_{\infty}\left(\bar{y}^{-1}I_{q},\bar{y}^{-3}I_{q},\ldots\right)^{*}{\bf h}_{2m}(y^{2}).$ (42)
@@
${\mathscr{G}}_{2m+1}^{(1)}(x,y)=$ ${\bf g}_{2m+1}^{*}(\bar{x}^{2})\left(I_{q},x^{-2}I_{q},\ldots\right)H^{(0)}_{\infty}\left(I_{q},\bar{y}^{-2}I_{q},\ldots\right)^{*}{\bf g}_{2m+1}(y^{2}),$ (43)
${\mathscr{G}}_{2m+1}^{(2)}(x,y)=$ $-{\bf g}_{2m+1}^{*}(\bar{x}^{2})\left(x^{-1}I_{q},x^{-3}I_{q},\ldots\right)H^{(1)}_{\infty}\left(\bar{y}^{-1}I_{q},\bar{y}^{-3}I_{q},\ldots\right)^{*}$
$\cdot{\bf g}_{2m+1}(y^{2}).$ (44)